Can you show 10 cent with 3 coins? Explain how you know
Answer by jim_thompson5910(35256)

No you cannot. Here are all the possible ways to have 3 coins (each coin is less than 10 cents)

penny, penny, penny = 1+1+1 = 3 cents

penny, penny, nickel = 1+1+5 = 7 cents

penny, nickel, nickel = 1+5+5 = 11 cents...which is NOT 10 cents since it's over 10 cents

nickel, nickel, nickel = 5+5+5 = 15 cents...which is NOT 10 cents since it's over 10 cents

So this shows us that it's not possible to make 10 cents with exactly 3 coins.